# Autonomous Driving in Unstructured Environments: How Far Have We Come?

### v3

@@ V Motion Control
V-A Challenges of Motion Control in Unstructured Environment
@@
V-B Proportional-Integral-Derivative (PID) Control
@@
V-C Preview Tracking Control
@@
V-D Sliding Mode Control
@@
V-E Model Predictive Control (MPC)
@@
Gibson et al. [42] uses a hybrid model with a specially-initialized LSTM for short-term dynamics prediction of an all-terrain vehicle, addressing error accumulation and scalability for sampling-based controllers. By limiting the LSTM to a fixed time horizon, it avoids long-term stability challenges. This flexible approach requires only odometry data for training. In [43], the system’s constraints are parallelized within the MPC, enabling efficient real-time planning and control. A low-level controller ensures the vehicle adapts to varying terrain without prior knowledge, ensuring safe, aggressive driving over challenging unstructured environments.
VI End-to-End Autonomous Driving
@@
TABLE XVII: End-to-end autonomous driving methods for unstructured environments.
@@
Muller et al. [237] Imitation Learning Behavior Cloning
Leininger et al. [44] Imitation Learning Behavior Cloning
Manderson et al. [45] Imitation Learning Behavior Cloning
Zhu et al. [46] Imitation Learning Inverse RL
Liang et al. [47] Online RL Policy-based
Weerakoon et al. [48] Online RL Actor-critic
@@
With the continuous advancement of deep learning technologies and the ongoing improvement of hardware computing capabilities, end-to-end autonomous driving methods have become a hot research topic for autonomous driving in unstructured environments [hassan2024terrainsense]. Depending on the different learning approaches of the network models, this method has gradually differentiated into two main research directions: imitation learning methods and online reinforcement learning methods, as shown in Table XVII.
VI-A Challenges of End-to-End Autonomous Driving in Unstructured Environment
@@ VI-B Imitation Learning
VI-B1 Behavior Cloning
Behavior cloning-based end-to-end autonomous driving methods train network models by minimizing the differences between human driving behavior strategies and those of ground unmanned platforms. This enables the platforms to simulate driving behavior strategies that closely resemble those of human drivers. Leininger et al. [44] propose a geometric-based framework for navigating uneven terrain without relying on pre-existing maps. This framework integrates a Sparse Gaussian Process (SGP) local map with a RRT* planner. The process begins with creating a high-resolution SGP local map, which provides an interpolated depiction of the robot’s immediate environment. This map captures key environmental features such as height variations, uncertainties, and slope characteristics. Next, a traversability map is generated from the SGP representation to guide the planning process. The RRT* planner then efficiently creates real-time navigation paths, avoiding untraversable terrain and directing the robot towards its goal. By combining SGP-based terrain interpretation with RRT* planning, ground robots can safely navigate areas with varying elevations and steep obstacles. The proposed approach’s effectiveness is validated through rigorous simulation testing, demonstrating notable improvements in safe and efficient navigation over existing methods.
@@ VI-B1 Behavior Cloning
VI-B2 Inverse Reinforcement Learning
@@
As shown in Fig. 9, Zhu et al. [46] introduce a method for off-road traversability analysis and trajectory planning using Deep Maximum Entropy Inverse Reinforcement Learning. Addressing the challenge of managing the exponential increase in state-space complexity while incorporating vehicle kinematics, the approach employs two convolutional neural networks: RL ConvNet and Svf ConvNet. These networks encode kinematic information into convolutional kernels, enabling efficient forward reinforcement learning. Experimental validation is conducted in off-road environments where scene maps are generated using 3D LiDAR data. Expert demonstrations consist of real vehicle driving trajectories in the scene or synthesized behaviors representing specific actions, such as traversing negative obstacles.
@@
VI-C Online Reinforcement Learning
@@
VI-C1 Value-based Methods
Value-based methods learn a value function through the interactive exploration process between the ground unmanned platform and the environment. The value function accurately assesses the expected return from executing a specific driving behavior strategy in a given state, allowing the platform to select the optimal driving behavior strategy by comparing the differences in expected returns. For example, Wolf et al. [241] propose an end-to-end autonomous driving framework based on Deep Q-Networks (DQN), which uses monocular images as input and learns human driving behavior strategies through a designed action-based reward function, ultimately achieving autonomous driving performance comparable to that of human drivers in a simulated environment. Min et al. [242] design a reward function that comprehensively considers factors such as speed, safety, and lane-changing frequency, and also utilized DQN to achieve autonomous driving in a highway simulation environment.
VI-C2 Policy-based Methods
Policy-based methods directly learn the probability distributions of various driving behavior strategies to be executed in a given state during the interactive exploration process, favoring the execution of the strategy with the highest probability. Liang et al. [47] introduce an outdoor navigation algorithm that ensures stable, efficient robot guidance. Using a multi-stage training pipeline with Proximal Policy Optimization (PPO), the method addresses drift, stability, and collision avoidance on complex terrains. It incorporates generalized environmental parameters and LiDAR data from a Unity simulator to bridge the sim-to-real gap. Evaluated on Clearpath Husky and Jackal robots, this approach effectively achieves robust navigation in both simulated and real-world settings.
VI-C3 Actor-critic Methods
@@
VII Datasets
@@
TABLE XVIII: Details of datasets for unstructured environments.
@@
Traversability Estimation TrailNet [49] RGB Camera - -
Vale [50] RGB Camera 600 4
CaT [51] RGB Camera 3,624 -
ORFD [52] RGB Camera, LiDAR 12,128 3
Verti-Wheelers [53] Stereo Camera 70,143 -
M3-GMN [54] LiDAR, RGB Camera, GPS/INS 1578 -
Semantic Segmentation DeepScene [55] RGB, NIR, Stereo 366 6
YCOR [56] RGB Camera 1,076 8
RUGD [57] RGB Camera 7,546 24
OFFSED [243] RGB, Stereo Camera 203 19
TAS500 [244] RGB Camera 640 24
RELLIS-3D [25] RGB, Stereo Camera 6,235 20
SynPhoRest [245] RGB Camera 1,974 6
BotanicGarden [246] RGB, Stereo Camera 1,181 27
GOOSE [58] RGB, NIR Camera 8,790 64
WildScenes [59] RGB Camera, LiDAR 12,148 15
Pedestrian Detection OPEDD [60] Stereo Camera 1,018 -
Dynamics Models TartanDrive [61] Stereo Camera 200,000 -
End-to-end Driving Rally Estonia [62] RGB Camera, LiDAR - -
Place Recognition Wild-Places [63] LiDAR 63,000 -
In recent years, data-driven deep learning algorithms have significantly advanced autonomous driving technology. Popular autonomous driving datasets such as KITTI [177], Waymo [247], and nuScenes [248] primarily focus on structured scenes. Recently, there has been increasing attention on publicly releasing autonomous driving datasets for unstructured environments, as shown in Fig. 10. As illustrated in Fig. 11 and Table XVIII, according to target tasks, these datasets can be categorized into traversability estimation, semantic segmentation, and other purposes.
VII-A Traversability Estimation
@@
VII-B Semantic Segmentation
@@ VII-C Other Datasets
VIII Future Perspectives
Future research efforts of autonomous driving in unstructured environments should concentrate on enhancing both modular autonomous driving and end-to-end autonomous driving, collecting more data, and boosting the safety and efficiency of autonomous driving systems. These endeavors will drive the advancement and application of autonomous driving technology in broader and more complex unstructured environments.
Offline Mapping. Future perspectives in offline mapping for autonomous driving in unstructured environments focus on creating detailed, high-resolution maps that account for diverse terrains and dynamic environmental changes. Advanced sensor technologies, such as LiDAR and photogrammetry [249, 250, 251], can capture the intricate features of landscapes, enabling the generation of comprehensive digital twins. Additionally, incorporating temporal data will allow these maps to adapt to seasonal and weather variations, ultimately enhancing navigation capabilities.
Pose Estimation. The challenge of pose estimation in unstructured environments necessitates the development of robust algorithms that leverage multiple sensor modalities, including cameras, LiDAR, and IMUs. Future advancements could delve into machine learning techniques that improve the understanding of complex spatial relationships and compensate for the ambiguity of features, thus enabling more accurate real-time localization in chaotic terrains.
Environmental Perception. In the realm of environmental perception, future innovations should focus on multi-modal sensor fusion that integrates visual, auditory, and tactile data to better identify and classify natural objects [252]. By employing advanced deep learning models, systems can more effectively discern the vast variety of elements in unstructured settings, leading to improved understanding of complex environments and the capability to adapt to new, unseen challenges [133].
Path Planning. Future path planning techniques for unstructured environments will likely utilize adaptive algorithms capable of dynamically adjusting routes based on real-time feedback from the environment. By incorporating reinforcement learning and predictive modeling, autonomous systems can effectively navigate disorderly scene structures and complex road conditions, optimizing safety and efficiency in uncertain terrains [253].
Motion Control. Advancements in motion control will prioritize the development of highly responsive systems that can handle the unpredictable nature of unstructured environments. Future perspectives include algorithms that enable vehicles to smoothly traverse varying surfaces and inclines while adapting to sudden changes in terrain, ultimately enhancing stability and control in challenging conditions.
End-to-End Autonomous Driving. The future of end-to-end autonomous driving in unstructured environments will emphasize the integration of advanced models that learn from vast datasets and real-world experiences. This holistic approach could enable systems to autonomously navigate complex scenes, recognizing and adapting to environmental variability without relying heavily on pre-defined rules or structures.
Dataset. Building robust datasets that encompass the diversity and complexity of unstructured environments is critical. Future efforts should focus on creating synthetic and real-world datasets that capture various scenarios, including seasonal changes and diverse terrains, to train machine learning models. Enhanced labeling techniques that address the ambiguity of semantic categories will also be essential for improving the accuracy and reliability of autonomous systems in these unpredictable settings [254].
IX Conclusion

### v4

@@ 5 Motion Control
5.1 Challenges of Motion Control in Unstructured Environment
@@
5.2 Proportional-Integral-Derivative (PID) Control
@@
5.3 Preview Tracking Control
@@
5.4 Sliding Mode Control
@@
5.5 Model Predictive Control (MPC)
@@
In complex and extreme off-road environments, vehicles frequently encounter uneven terrain, mixed soft and hard soils, and unpredictable slopes or obstacles, posing significant challenges to traditional fixed-parameter controllers. Adaptive control has therefore become a key research focus. [143] present an online adaptive Model Predictive Control (MPC) framework for off-road vehicles operating on deformable terrains, incorporating a physics-informed, online-adaptable tire model. Simulation results demonstrate that online adaptation significantly improves speed and path-tracking performance when discrepancies exist between assumed and actual terrain parameters. [51] use a hybrid model with a specially-initialized LSTM for short-term dynamics prediction of an all-terrain vehicle, addressing error accumulation and scalability for sampling-based controllers. By limiting the LSTM to a fixed time horizon, it avoids long-term stability challenges. This flexible approach requires only odometry data for training. In [63], the system’s constraints are parallelized within the MPC, enabling efficient real-time planning and control. A low-level controller ensures the vehicle adapts to varying terrain without prior knowledge, ensuring safe, aggressive driving over challenging unstructured environments.
5.6 Reinforcement Learning-based Control
Another class of reinforcement learning (RL)-based controllers focuses on learning strategies to compensate for slip and wheel-terrain interactions on soft surfaces, enabling vehicles to adapt to unknown or extreme terrains. For example, [61] proposes an Actor-Critic reinforcement learning Compensated Model Predictive Controller (AC2MPC) for high-speed off-road driving on deformable terrains, integrating deep RL with MPC to handle unmodeled tire-terrain dynamics. Simulation results demonstrate that the approach outperforms standalone model-based or learning-based controllers across multiple unknown terrains, achieving more accurate speed tracking, faster convergence, and strong generalization with limited training data. These methods highlight the significant potential of adaptive control to enhance vehicle stability and performance in complex terrains, though challenges remain in real-time computational efficiency, perception uncertainty, and cross-terrain generalization.
5.7 Model Predictive Path Integral (MPPI)
In off-road autonomous driving, vehicles frequently encounter complex terrains, dynamic obstacles, and highly nonlinear vehicle dynamics, posing significant challenges for traditional optimization- or linearization-based controllers. Recently, Model Predictive Path Integral (MPPI) control, a sampling-based nonlinear optimal control method, has gained widespread attention in off-road environments. MPPI performs stochastic trajectory sampling and cost-weighted averaging directly in the trajectory space, enabling real-time handling of nonlinear, constrained problems without relying on precise vehicle models. This makes it well-suited for rugged terrains, mixed soft and hard soils, and nonlinear tire-terrain interactions.
Studies have demonstrated MPPI’s advantages in off-road autonomous navigation. For example, [216] proposed an MPPI-based off-road planning method that samples trajectories in high-dimensional state space, achieving robust obstacle avoidance and path tracking in unknown or dynamic environments. [227] integrated MPPI with vehicle dynamics constraints and terrain information for high-speed off-road driving, showing strong adaptation to soft ground slip and tire-terrain interactions. Furthermore, MPPI’s parallelizable computation allows near real-time operation on GPUs or embedded edge platforms, making it feasible for practical deployment. Despite these strengths, MPPI still faces challenges in sampling efficiency, cost function design, and coordinated planning for multi-vehicle systems. Overall, MPPI-based controllers provide an effective framework for autonomous navigation in complex off-road environments and offer important directions for adaptive and high-robustness control research.
6 End-to-End Autonomous Driving
@@
Table 17: End-to-end autonomous driving methods for unstructured environments.
@@
[135] Imitation Learning Behavior Cloning
[100] Imitation Learning Behavior Cloning
[120] Imitation Learning Behavior Cloning
[256] Imitation Learning Inverse RL
[108] Online RL Policy-based
[209] Online RL Actor-critic
@@
With the continuous advancement of deep learning technologies and the ongoing improvement of hardware computing capabilities, end-to-end autonomous driving methods have become a hot research topic for autonomous driving in unstructured environments [64]. Depending on the different learning approaches of the network models, this method has gradually differentiated into two main research directions: imitation learning methods and online reinforcement learning methods, as shown in Table 17.
6.1 Challenges of End-to-End Autonomous Driving in Unstructured Environment
@@ 6.2 Imitation Learning
6.2.1 Behavior Cloning
Behavior cloning-based end-to-end autonomous driving methods train network models by minimizing the differences between human driving behavior strategies and those of ground unmanned platforms. This enables the platforms to simulate driving behavior strategies that closely resemble those of human drivers. [100] propose a geometric-based framework for navigating uneven terrain without relying on pre-existing maps. This framework integrates a Sparse Gaussian Process (SGP) local map with a RRT* planner. The process begins with creating a high-resolution SGP local map, which provides an interpolated depiction of the robot’s immediate environment. This map captures key environmental features such as height variations, uncertainties, and slope characteristics. Next, a traversability map is generated from the SGP representation to guide the planning process. The RRT* planner then efficiently creates real-time navigation paths, avoiding untraversable terrain and directing the robot towards its goal. By combining SGP-based terrain interpretation with RRT* planning, ground robots can safely navigate areas with varying elevations and steep obstacles. The proposed approach’s effectiveness is validated through rigorous simulation testing, demonstrating notable improvements in safe and efficient navigation over existing methods.
@@ 6.2.1 Behavior Cloning
6.2.2 Inverse Reinforcement Learning
@@
As shown in Fig. 8, [256] introduce a method for off-road traversability analysis and trajectory planning using Deep Maximum Entropy Inverse Reinforcement Learning. Addressing the challenge of managing the exponential increase in state-space complexity while incorporating vehicle kinematics, the approach employs two convolutional neural networks: RL ConvNet and Svf ConvNet. These networks encode kinematic information into convolutional kernels, enabling efficient forward reinforcement learning. Experimental validation is conducted in off-road environments where scene maps are generated using 3D LiDAR data. Expert demonstrations consist of real vehicle driving trajectories in the scene or synthesized behaviors representing specific actions, such as traversing negative obstacles.
@@
6.3 Online Reinforcement Learning
@@
6.3.1 Value-based Methods
Value-based methods learn a value function through the interactive exploration process between the ground unmanned platform and the environment. The value function accurately assesses the expected return from executing a specific driving behavior strategy in a given state, allowing the platform to select the optimal driving behavior strategy by comparing the differences in expected returns. For example, [218] propose an end-to-end autonomous driving framework based on Deep Q-Networks (DQN), which uses monocular images as input and learns human driving behavior strategies through a designed action-based reward function, ultimately achieving autonomous driving performance comparable to that of human drivers in a simulated environment. [130] design a reward function that comprehensively considers factors such as speed, safety, and lane-changing frequency, and also utilized DQN to achieve autonomous driving in a highway simulation environment.
6.3.2 Policy-based Methods
Policy-based methods directly learn the probability distributions of various driving behavior strategies to be executed in a given state during the interactive exploration process, favoring the execution of the strategy with the highest probability. [108] introduce an outdoor navigation algorithm that ensures stable, efficient robot guidance. Using a multi-stage training pipeline with Proximal Policy Optimization (PPO), the method addresses drift, stability, and collision avoidance on complex terrains. It incorporates generalized environmental parameters and LiDAR data from a Unity simulator to bridge the sim-to-real gap. Evaluated on Clearpath Husky and Jackal robots, this approach effectively achieves robust navigation in both simulated and real-world settings.
6.3.3 Actor-critic Methods
@@
7 Multi-Agent Collaborative Systems
In recent years, Multi-Agent Collaborative Systems (MACS) have emerged as a research hotspot in the field of off-road autonomous driving [197]. Compared with single-vehicle autonomous navigation, multi-vehicle (or multi-robot) collaborative systems achieve more efficient environmental perception, path planning, and decision-making through information sharing and task allocation, making them particularly suitable for unstructured scenarios such as disaster search and rescue, mining transportation, and agricultural operations. In this context, [29] conduct a comprehensive survey on multi-agent reinforcement learning methods for cooperative driving, analyzing key challenges such as communication constraints and distributed policy optimization in collaborative perception and decision-making. [243] propose a distributed multi-robot cooperative exploration framework based on Graph Neural Networks (GNNs), enabling robots to achieve globally efficient mapping in complex and rugged terrains. The CTU-CRAS-NORLAB team developed a multi-robot exploration framework for ground robots, focusing on planning, traversability estimation, and decentralized coordination in large-scale, communication-limited underground environments during the 2021 DARPA Subterranean Challenge [8]. The framework combines dense local mapping with a shareable sparse topometrical map, enabling fully autonomous exploration and effective multi-robot coordination, as demonstrated in both Virtual and Systems tracks of the competition.
8 Datasets
@@
In recent years, data-driven deep learning algorithms have significantly advanced autonomous driving technology. Popular autonomous driving datasets such as KITTI [50], Waymo [182], and nuScenes [13] primarily focus on structured scenes. Recently, there has been increasing attention on publicly releasing autonomous driving datasets for unstructured environments, as shown in Fig. 9. As illustrated in Fig. 10 and Table 18, according to target tasks, these datasets can be categorized into traversability estimation, semantic segmentation, and other purposes.
Table 18: Details of datasets for unstructured environments.
@@
Traversability Estimation TrailNet RGB Camera - -
Vale RGB Camera 600 4
CaT RGB Camera 3,624 -
ORFD RGB Camera, LiDAR 12,128 3
Verti-Wheelers Stereo Camera 70,143 -
M3-GMN LiDAR, RGB Camera, GPS/INS 1578 -
ORAD-3D LiDAR, RGB Camera 57,808 -
Semantic Segmentation DeepScene RGB, NIR, Stereo 366 6
YCOR RGB Camera 1,076 8
RUGD RGB Camera 7,546 24
TAS500 RGB Camera 640 24
RELLIS-3D RGB, Stereo Camera 6,235 20
BotanicGarden RGB, Stereo Camera 1,181 27
GOOSE RGB, NIR Camera 8,790 64
WildScenes RGB Camera, LiDAR 12,148 15
ORAD-3D LiDAR, RGB Camera 57,808 8
Pedestrian Detection OPEDD Stereo Camera 1,018 -
Dynamics Models TartanDrive Stereo Camera 200,000 -
End-to-end Driving Rally Estonia RGB Camera, LiDAR - -
ORAD-3D LiDAR, RGB Camera 57,808 -
Place Recognition Wild-Places LiDAR 63,000 -
Path Planning Trajectory_Generate LiDAR, RGB-D Camera, GNSS 20,859 -
8.1 Traversability Estimation
@@
ORAD-3D [126] establishes a comprehensive suite of benchmark evaluations spanning five fundamental tasks: 2D free-space detection, 3D occupancy prediction, rough GPS-guided path planning, vision-language model-driven autonomous driving, and world model for off-road environments. Together, the dataset and benchmarks provide a unified and robust resource for advancing perception and planning in challenging off-road scenarios.
8.2 Semantic Segmentation
@@ 8.3 Other Datasets
Trajectory_Generate [190] uses an RGB-D camera combined with an IMU to collect data for image-based end-to-end path planning. The data was collected from three different locations: a university greenhouse, a university park, and a nearby natural trail.
9 Future Perspectives
To advance the development of autonomous driving technology in unstructured environments, future research should focus on the following core areas:
High-Precision Dynamic Map Construction. Develop high-resolution offline maps that capture terrain details and dynamic environmental changes using LiDAR and photogrammetry [254, 211, 212]. Emphasize spatiotemporal modeling to address seasonal and weather variations by integrating time-series data (e.g., monthly point clouds) and employing techniques such as 4D Gaussian processes or spatiotemporal CNNs. A digital twin system with dynamic updates can provide a high-precision, temporally-aware knowledge base to support navigation, path planning, and decision-making in complex, changing terrains.
Multimodal Pose Estimation System. To address the challenge of ambiguous terrain features, new algorithms for camera-LiDAR-IMU multi-sensor fusion need to be developed. Using deep learning methods to analyze complex spatial relationships, a real-time positioning framework based on probabilistic reasoning should be created to solve pose estimation issues in feature-deficient scenarios, thereby improving system robustness in dynamic environments.
Environmental Perception Enhancement. Build a multimodal perception network combining vision, LiDAR, and RADAR  [103], leveraging graph neural networks and Transformer architectures to achieve fine-grained semantic understanding of natural objects. The focus should be on enhancing open-set recognition in unstructured environments and developing explainable perception models [168] to enable adaptive classification and risk prediction for unknown obstacles.
Intelligent Path Planning and Decision Making. Develop a dynamic path planning framework based on reinforcement learning, integrating real-time terrain scanning and predictive environmental modeling. By designing cost functions that account for risk perception, a multi-objective decision system [206] should be created to optimize both safety and energy consumption, enabling collaborative optimization of global path planning and local obstacle avoidance in complex terrains.
Nonlinear Motion Control Architecture. Develop an adaptive controller with terrain prediction capabilities, and create a hybrid control strategy that combines model predictive control and learning methods. The focus should be on addressing dynamic coupling issues in extreme conditions, such as steep slopes and soft ground, while establishing a slip compensation mechanism that considers tire-ground interaction characteristics, improving motion stability in rapidly changing terrain.
End-to-End System Integration and Innovation. Create a neural radiance field simulation environment for off-road scenarios and develop an end-to-end driving architecture based on world models. By combining imitation learning and reinforcement learning, the system should be capable of making end-to-end decisions from raw sensor inputs to control outputs, minimizing reliance on manually defined rules.
Multidimensional Dataset Construction. Existing unstructured off-road datasets remain limited in scale and diversity, highlighting the need to expand their richness in future work. Current datasets also provide little analysis of semantic ambiguities, such as distinguishing between passable grass and dense tall vegetation, which should be addressed to improve perception models. Moreover, most existing datasets focus on static environments, while dynamic elements—such as swaying branches or flowing water—are rarely captured, limiting their applicability for training dynamic obstacle avoidance algorithms. Future efforts should therefore aim to collect datasets that include both dynamic elements and terrain–vehicle interactions.
Extreme Environmental Conditions. Future research should systematically evaluate the performance of off-road autonomous driving algorithms under extreme environmental conditions, such as heavy rain, snow, or dynamic obstacles like wildlife and falling rocks. For instance, the impact of LiDAR point cloud noise caused by rain on traversability estimation accuracy remains largely unexplored. Investigating sensor robustness, perception uncertainty, and adaptive planning strategies under such conditions will be critical for achieving reliable and safe off-road autonomous navigation in real-world scenarios.
Hardware Heterogeneity. Future studies should investigate the impact of heterogeneous hardware configurations on the performance of off-road autonomous driving algorithms. For example, differences between low-resolution (16-beam) and high-resolution (128-beam) LiDAR, or between consumer-grade and high-end IMUs, can significantly affect perception accuracy, localization, and control. Quantifying these effects will help researchers and practitioners select appropriate hardware for specific scenarios and optimize system performance under resource constraints.
Edge Computing Integration. Off-road autonomous driving in unstructured environments faces challenges such as weak networks, high latency, and large sensor data [115]. Traditional cloud computing often fails to meet real-time demands. Future work should leverage edge computing for local and collaborative sensor processing, including LiDAR compression, semantic extraction, and dynamic object recognition. Key directions include task allocation, computation offloading, robustness under limited connectivity, and scalable multi-vehicle coordination. Progress in these areas is essential for fully autonomous, resilient off-road operation.
Explainable Research for Off-Road AD. Future work should develop explainable perception and decision-making methods for unstructured off-road environments, where complex terrain, dynamic obstacles, and sensor noise challenge autonomy. Techniques that highlight which LiDAR points, camera features, or fused sensor data drive obstacle detection, traversability estimation, and planning decisions are essential to enable real-time validation, build operator trust, and ensure safety in critical applications such as military missions and disaster response.
10 Conclusion